2.3 Solve the following linear programming graphically [5]

Minimize 𝑧 = 3π‘₯ + 9𝑦

Subject to the constraints: π‘₯ + 3𝑦 β‰₯ 6

π‘₯ + 𝑦 ≀ 10

π‘₯ ≀ 𝑦

π‘₯ β‰₯ 0; 𝑦 β‰₯ 0


1
Expert's answer
2021-10-12T05:09:25-0400

Minimize "\ud835\udc67 = 3\ud835\udc65 + 9\ud835\udc66"

subject to the constraints


"\begin{matrix}
 x+3y\geq6 \\
x+y \leq 10 \\
x \leq y \\
x\geq0, y\geq0
\end{matrix}"

Find the point(s) of intersection


"y=-\dfrac{1}{3}x+2"

"y=-x+10"

"x=0:"


"y=-\dfrac{1}{3}(0)+2, Point\ A(0,2)"

"y=-0+10, Point\ B(0,10)"

"-\dfrac{1}{3}x+2=x"

"\dfrac{4}{3}x=2"

"x=1.5, y=1.5,Point\ D(1.5,1.5)"


"-x+10=x"

"2x=10"

"x=5, y=5,Point\ C(5,5)"


Point "A(0,2):\ud835\udc67(0,2) =3(0) + 9(2)=18"


Point "B(0,10):\ud835\udc67(0,10) = 3(0) +9(10)=90"


Point "C(5,5):\ud835\udc67(5,5) = 3(5) +9(5)=60"


Point "D(1.5,1.5):\ud835\udc67(1.5,1.5) = 3(1.5) +9(1.5)=18"


The function "z" has a minimum with value of "18" at "(0,2)" and at "1.5,1.5)."